OpenStudy (anonymous):

Given: F(x) = x + 2 and G(x) = 3x + 5 (F + G) (x) =

OpenStudy (lalaly):

(F+G)(x)=F(x)+G(x)

Parth (parthkohli):

Just add the two functions together.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (f + g)(x) = f(x) + g(x) = x + 2 + 3x + 5 }\)